OpenStudy (anonymous):

express log2 64-log2 4 as a simple logarithm

OpenStudy (anonymous):

Since they both have base 2, u can combine them. Recall, subtraction of logs means division. log2 64 - log2 4 = log2 (64/4) = log2 (16)

OpenStudy (anonymous):

what math1234 said. and since \[2^4=16\] the answer is 4